OpenStudy (anonymous):

Need Help! Factor each polynomial completely. Write any repeated factor in exponential form, then name all zeroes and their multiplicity. P(X)= (X ^2 -10X + 25) (X^2 + 4X - 45) (X+9)

OpenStudy (anonymous):

ok so here we can start to simplify the polynomials. simplifying first poly: x^2 -10x + 25 we get (x-5)(x-5) simplifying the second bracket: x^2 + 4x - 45 we get (x+9)(x-5) so now writing P(x) = (x-5)(x-5)(x+9)(x-5)(x+9) = (x+5)^3 * (x+9)^2 soo from here ... x+9 = 0 so x = -9 x+5 = 0 so x = -5 Hence x = -5 and -9 are two answers for the solution